# A Classification Theorem for Steady Euler Flows

### v1

@@ 3 Structure of the set of critical points
Let $\Gamma$ be a curve of points in $\mathcal{C}.$ Suppose that for some $x_{*}\in\Gamma$, $d_{\psi}(x_{*})=2.$ Then, $d_{\psi}(x)=2$ for every $x$ in the connected component of $\{\nabla\psi=0\}$ containing $\Gamma.$
@@
By Proposition 2.6, for a non-isolated critical point $x$, we have that $D^{2}\psi(x)=0$ is equivalent to $\Delta\psi=0$; moreover, $\Delta\psi$ is constant along any curve. This is because, if we take any analytic curve emanating from $x,$ and denote by $\tau$ and $n$ the tangent and normal vectors to the curve at $x,$ we have that $\partial_{\tau n}\psi(x)=\partial_{\tau\tau}\psi(x)=0.$
@@
The points for which $d_{\psi}(x)=2$ consist of a collection of isolated points or isolated analytic closed loops in $\Omega$.
@@
Let $\Gamma$ be an analytic curve of points of $\psi$ of degree $d\geq 3$. Then, for all $x\in\Gamma,$
@@
Fix $x\in\Gamma$. By the rotational invariance of $\Delta,$ we may assume that the tangent vector to $\Gamma$ at $x$ is $(1,0).$ Now, for each $y\in\Gamma$ we have $D^{d-1}\psi(y)=0$, from which we conclude that $D^{d-1}\partial_{1}\psi(x)=0$. It thus follows that $D^{d}\psi(x)\not=0$ is equivalent to $\partial_{2}^{d}\psi(x)\not=0.$ This, in turn, is equivalent to $\partial_{2}^{d-2}\Delta\psi(x)\not=0.$
@@
3.1 The Inner-Most Loop
Since $\mathcal{C}(\psi)$ contains curves, $\Omega$ must be partitioned by the level set $\{\nabla\psi=0\}$ (this follows from the above discussion and Proposition 2.6). In particular, we can find at least one curve of critical points that is either a closed loop in $\Omega$ or starts and ends at $\partial\Omega.$ Let $\Gamma$ be one such curve with the property that the area it encloses contains no curves of critical points. Let us call this enclosed domain $\Omega_{\Gamma},$ which is necessarily simply connected. All of the analysis below will be restricted to $\Omega_{\Gamma}.$ Since $\Omega_{\Gamma}$ only contains isolated critical points of $\psi,$ we have that
$\Delta\psi=F(\psi)\text{ in $\overline{\Omega_{\Gamma}},$ }$ (3.1)
@@ 3.1 The Inner-Most Loop
$\psi|_{\partial\Omega_{\Gamma}}=0.$
We will now show that we have, in fact, that $\nabla\psi|_{\partial\Omega_{\Gamma}}=0,$ which is automatic when $\Gamma$ is a loop.
@@
If $\Gamma$ touches $\partial\Omega,$ then the portion of $\partial\Omega$ between the two endpoints of $\Gamma$ lies in $\mathcal{C}(\psi).$
@@
By Corollary 3.2, if $\Gamma$ touches $\partial\Omega$, then $d_{\psi}(x)\geq 3$ for all $x\in\Gamma.$ Moreover, $\psi$ must then vanish on $\partial\Omega.$ Now, take some $x\in\partial\Omega.$ If $\nabla\psi(x)\not=0,$ then $F$ from (3.1) must be analytic in a neighborhood of $0$ In addition, the condition that $d_{\psi}(x)\geq 3$ for all $x\in\Gamma$ shows $F(0)=0$. This, however, contradicts Lemma 3.3, since $F$ being analytic with $F(0)=0$ would imply that $\Delta\psi=F(\psi)$ vanishes to the same order (or higher) than $\psi$ on $\Gamma.$
@@
One of the following two can hold
  • Case 1: $\partial\Omega_{\Gamma}$ is analytic and contains only critical points of degree 2.
  • Case 2: $\partial\Omega_{\Gamma}$ is a closed loop containing critical points of degree larger than 2.
Our goal will now be to show that, in both cases, $\psi$ is necessarily a radial function. We will first deal with Case 2, which will be solved essentially by applying existing results from the literature. Case 1 will be dealt with in the following section using the moving plane method.
3.2 Rigidity in Case 2
In the previous subsection, we reduced the problem into studying an semi-linear elliptic equation of Serrin type [39]. More precisely, we found a simply connected domain $\Omega_{\Gamma}$ for which $\nabla\psi$ vanishes on $\partial\Omega_{\Gamma}$ and for which $\Omega_{\Gamma}$ contains only isolated critical points of $\psi$. As a consequence of Lemma 1.4, we deduced that $\psi$ solves (3.1) in $\Omega_{\Gamma}$. In this section, we consider Case 2 of Corollary 3.6, where all the points of $\partial\Omega_{\Gamma}$ are critical points with degree greater than or equal to $3$. Our goal will be to establish the following proposition.
@@ 3.2 Rigidity in Case 2
Let $\Omega_{\Gamma}$ be a simply connected domain and $\psi$ be an analytic function in $\overline{\Omega_{\Gamma}}$ such that there is no non-isolated critical points of $\psi$ in $\Omega_{\Gamma}\cap\{\psi=0\}$. Now we further assume there is a continuous function $F$, such that $\psi$ satisfies the following over-determined elliptic system.
$\Delta\psi(x)=F(\psi)(x),\text{ for $x\in\Omega_{\Gamma}$,}$ (3.2)
$\psi(x)=\nabla\psi(x)=\nabla^{2}\psi(x)=0,\text{ for all $x\in\partial\Omega_{\Gamma}$},$
@@
We first recall the following result based on Steiner symmetrization method.
@@
Let $\Omega$ be an open bounded set and $\psi$ a non-negative $C^{2}$ solution to
[…]
Then $F$ being continuous implies that $\psi$ is locally radial that is:
@@
where $A=\cup_{i}A_{i}$, $\mathcal{C}=\{\nabla\psi=0\}$ and $A_{k}$ are disjoint open annuli or balls and $\psi$ is radially symmetric in $A_{k}$.
@@
Given Proposition 3.8, all we need to do is show that $\psi$ is single signed on $\Omega_{\Gamma}.$ Towards a contradiction, assume that $\psi$ takes both positive and negative values in $\Omega_{\Gamma}.$ Simply by continuity, it follows that the set of zeros of $\psi$ in $\Omega_{\Gamma}$ is infinite. It follows, by our choice of $\Gamma,$ that at least one of those zeros is not a critical point of $\psi$ (since the critical set in $\Omega_{\Gamma}\cap\{\psi=0\}$ is discrete). It follows that $F$ is analytic in a neighborhood of zero. This contradicts Lemma 3.3 (just like in the proof of Lemma 3.5). ∎
4 Moving Plane Argument
We now move on to prove the rigidity in Case 1 of Corollary 3.6. In particular, we may assume that there exists a simply connected domain $\Omega_{\Gamma}$ with $\partial\Omega_{\Gamma}$ analytic with $\psi$ satisfying
$\Delta\psi(x)=F(\psi)(x),\text{for $x\in\Omega_{\Gamma}$,}$ (4.1)
@@ 4 Moving Plane Argument
$\partial_{nn}\psi|_{\partial\Omega_{\Gamma}}=1.$
An important consequence is that $\psi$ is increasing in the direction normal to the boundary, in a tubular neighborhood of the boundary. In the case where $F$ is $C^{1}$ and $\psi$ is non-negative, we have the following rigidity result due to Serrin [39] (see also the nice review [36]).
Theorem 4.1 (Serrin, 1971).
Assume that $\psi$ satisfies (4.1) and that $\psi>0$ in $\Omega_{\Gamma},$ while $F$ is Lipschitz continuous. Then, $\Omega_{\Gamma}$ is a ball and $\psi$ is radial.
@@
It is either that
  • $F(s)$ is not analytic at $s=0,$ in which case $\psi$ is non-negative on $\Omega_{\Gamma}.$ This is called the unextendable case.
@@
4.1 The Unextendable Case
This is the case where $F(s)$ is not analytic in $s$ at 00. As mentioned in Lemma 4.2, $\psi$ must be non-negative in $\Omega_{\Gamma}$, since otherwise there would be infinitely many points where $\psi=0$ in $\Omega_{\Gamma}$. Without loss of generality, by scaling and multiplication by a constant factor, we assume that $\max_{\Omega_{\Gamma}}\psi=1$. Analogously as in the proof of Proposition 3.7, the condition that $\psi$ does not have a non-isolated critical point in $\Omega_{\Gamma}$ implies that $F$ is analytic in $(0,1)$. Our goal in this section is to prove the following proposition:
@@ 4.1 The Unextendable Case
Let $\Omega_{\Gamma}$ be a simply connected analytic domain, $\psi$ be a non-negative analytic function in $\overline{\Omega_{\Gamma}}$. Assume $F\in C([0,1])$ is analytic on $(0,1).$ If $F$ is not analytic at 00 and $\psi$ satisfies the over-determined elliptic equation (4.1), then $\Omega_{\Gamma}$ is a ball and $\psi$ is radial.
@@
Without a loss of generality, we may assume $F(0)=1,$ since the case $F(0)=0$ is already covered by Proposition 3.7.
For the proof of Proposition 4.3, it suffices to prove that $\psi$ is even symmetric to all directions $e\in\mathbb{S}^{1}$. Without loss of generality, by rotation, $e=(1,0)$. We will use the moving plane method to show that $\psi$ is even symmetric to $(1,0)$. We start with a series of lemmas concerning the regularity of $F$ and the critical set of $\psi$. First, either $F$ is analytic at $1$ or $F$ is not smooth and admits a Puiseux expression near $1$.
@@
When the maxima of $\psi$ are critical points of degree $2k_{0},$ for some $k_{0}\geq 1,$ $F(s)$ admits the following convergent expansion.
$F(s)=\sum_{k=k_{0}-1}a_{k}(1-s)^{\frac{k}{k_{0}}},\text{ for $0\leq 1-s\ll 1$,with $a_{k}$ real and $a_{k_{0}-1}<0$. }$ (4.2)
@@
Fix a point $x^{0}\in\Omega_{\Gamma}$ such that $\psi(x^{0})=1$, as $\psi$ attains the global maximum at $x^{0}$, we have $d_{\psi}(x^{0})$ is an even number. In the case $d_{\psi}(x^{0})=2$, without loss of generality, by a suitable rotation and translation, we assume $x^{0}=0$ and $\partial_{11}\psi(0)<0$. As $\partial_{1}\psi(0)=0$ and $\partial_{11}\psi(0)<0$, for $0\leq 1-s\ll 1$, the function $\psi(t,0)$ near 00 has exactly two inverses $f_{+}(s)$ and $f_{-}(s)$, with $f_{+}(s)\geq 0$ and $f_{-}(s)\leq 0$. We show that $f_{+}$ and $f_{-}$ admit a Puisuex expression near $1$. In particular, we will show that there is an analytic function $G$, such that $f_{+}(s)=G(\sqrt{1-s})$ and $f_{-}(s)=G(-\sqrt{1-s})$. Indeed given the Taylor expression of $\psi(t,0)=1+a_{2}t^{2}+\sum_{n\geq 3}a_{n}t^{n}$ with $a_{2}<0$, we have
$\psi(f_{\pm},0)=s\Leftrightarrow f_{\pm}^{2}\left(1+\sum_{k\geq 3}\frac{a_{3}}{a_{2}}f_{\pm}^{k-2}\right)=\frac{s-1}{a_{2}}\Leftrightarrow f_{\pm}\sqrt{1+\sum_{k\geq 3}\frac{a_{3}}{a_{2}}f_{\pm}^{k-2}}=\pm\sqrt{\frac{s-1}{a_{2}}}.$ (4.3)
Now the function $s\sqrt{1+\sum_{k\geq 3}\frac{a_{3}}{a_{2}}s^{k-2}}$ is an analytic function with a non-vanishing derivative at 00, hence it admits an analytic inverse $H$ and we can correspondingly choose $G(s)=H(\frac{s}{\sqrt{-a_{2}}})$. With the aid of $f_{+}$, $f_{-}$ and $G$ in the claim, for $0\leq 1-s\ll 1$, we have
@@
The fact that $\Delta\psi(G(\sqrt{1-s}),0)$ agrees with $\Delta\psi(G(-\sqrt{1-s}),0)$ shows that $F$ is analytic near $1$.
Now in the case where $d_{\psi}(x^{0})=2k_{0}\geq 4$, again assume that $x^{0}=0$ and $\partial_{1}^{2k_{0}}\psi<0$, for $0\leq 1-s\ll 1$, the function $\psi(t,0)$ near 00 has exactly two inverses $f_{+}(s)$, $f_{-}(s)$. Similarly as in the case where $d_{\psi}(x^{0})=2$, we can show that $F$ near $1$ has the Puiseux expression.
$F(s)=\sum_{k=k_{0}-1}a_{k}(1-s)^{\frac{k}{k_{0}}},\text{with $a_{k}$ real, $a_{k_{0}-1}<0$, for $0\leq 1-s\ll 1$. }$ (4.4)
Analogously, by taking a normal line segment through a boundary point $x\in\partial\Omega_{\Gamma}$, we have $F(s)$ is analytic in the variable $\sqrt{s}$ near 00. Together with the condition that $F$ is not analytic near 00, we have the following Puiseux expression of $F$ near 00.
@@
$F$ near 00 admits a convergent Puiseux series expression:
$F(s)=\sum_{k=0}a_{k}s^{\frac{k}{2}},\text{for $0\leq s\ll 1$ , with $a_{k}$ real, $a_{0}=1$ and $a_{k_{0}}\neq 0$ for an odd number $k_{0}$.}$ (4.5)
Moreover, by arguing similarly as in the case $d_{\psi}(x^{0})=2$ in the proof of Lemma 4.5, the condition $F$ is not analytic at 00 implies that $\psi$ is positive in $\Omega_{\Gamma}$. Together with fact
@@
we can show $\psi$ has a uniform lower bound in $\Omega_{\Gamma}$.

### v2

@@ 3. Structure of the Set of Critical Points
Let $\Gamma_{0}$ be a curve of points in $\mathcal{C}.$ Suppose that for some $x_{*}\in\Gamma_{0}$, $d_{\psi}(x_{*})=2.$ Then, $d_{\psi}(x)=2$ for every $x$ in the connected component of $\{\nabla\psi=0\}$ containing $\Gamma_{0}.$
@@
We first choose $\delta$ small, such that the Hessian matrix $D^{2}\psi(z)\neq 0$ for all $z\in B_{\delta}(x_{*})$. Based on Proposition 2.3, then there exists a $y\in B_{\delta}(x_{*})\cap\Gamma_{0}$ such that $\Gamma_{0}$ near $y$ is an analytic curve. Now pick $t$ and $n$ to be the tangent and normal direction of $\Gamma_{0}$ at $y$, respectively. The condition that $\Gamma_{0}$ consists of critical points of $\psi$ implies that $\partial_{tt}\psi(y)=\partial_{tn}\psi(y)=0$. Now, as $D^{2}\psi(y)\neq 0$, we have $\Delta\psi(y)=\partial_{tt}\psi(y)+\partial_{nn}\psi(y)\neq 0$. As $\{\psi,\Delta\psi\}=0$, we have that $\Delta\psi$ is constant along each connected component of $\{\nabla\psi=0\}$. In particular, we have $\Delta\psi(x)\neq 0$ and thus $d_{\psi}(x)=2$ for every $x$ in the connected component of $\{\nabla\psi=0\}$ containing $x_{*}$. This finishes the proof of the lemma. ∎
@@
The points in $\Omega$ for which $d_{\psi}(x)=2$ consist of a collection of isolated points or isolated analytic closed loops in $\Omega$.
Proof.
Recall that we are assuming that $\psi$ is analytic on $\bar{\Omega}.$ It follows then from Lemma 2.4 that if a curve $\Gamma_{0}$ of critical points of degree $2$ from the interior were to hit $\partial\Omega$, then the degree at the supposed intersection point, $x\in\partial\Omega$, would have to be at least $3$. This would contradict the constancy of $\Delta\psi$ along $\Gamma_{0},$ as in Lemma 3.1. ∎
@@
Let $d\geq 3$ and assume that $\Gamma_{0}$ be an analytic curve for which $d_{\psi}(x)=d,$ for all $x\in\Gamma_{0}$. Then, for all $x\in\Gamma_{0},$
@@
It is clear that $D^{i}\Delta\psi(x)=0$ for all $x\in\Gamma_{0}$ and $i\leq d-3$. Hence it suffices to check $D^{d-2}\Delta\psi(x)\not=0$ for all $x\in\Gamma_{0}$. Fix $x^{0}\in\Gamma_{0}$. By the rotational invariance of $\Delta,$ we may assume that the tangent vector to $\Gamma_{0}$ at $x^{0}$ is $(1,0).$ Now, for each $y\in\Gamma_{0}$ we have $D^{d-1}\psi(y)=0$. It follows that $D^{d-1}\partial_{1}\psi(x^{0})=0$. Now $D^{d}\psi(x^{0})\not=0$ implies that $\partial_{2}^{d}\psi(x^{0})\not=0$. Hence, one has $D^{d-2}\Delta\psi(x^{0})\not=0$. This concludes the proof. ∎
@@
3.1. The Innermost Loop
Since $\mathcal{C}(\psi)$ contains curves, $\Omega$ must be partitioned by the level set $\{\nabla\psi=0\}$ (this follows from the above discussions and Proposition 2.3). In particular, we can find at least one curve of critical points that is either a closed loop in $\Omega$ or starts and ends at $\partial\Omega$. Note that in the case where the curve starts and ends at $\partial\Omega$, it must touch the boundary at two points (the two points could be the same). Together with a boundary portion between these two points, the curve would still enclose a closed simply connected region. Now, let $\Gamma$ be a curve with the property that the region it encloses (possibly with the aid of a portion of the boundary) contains no curves of critical points. Let us call this enclosed domain $\Omega_{\Gamma},$ which is necessarily simply connected.
All of the analysis below will be restricted to $\Omega_{\Gamma}.$ Since the critical points of $\psi,$ contained in $\Omega_{\Gamma}$ must be isolated, we have that
(3.1) $\Delta\psi=F(\psi)\text{ in $\overline{\Omega_{\Gamma}},$ }$
@@ 3.1. The Innermost Loop
$\psi=0\quad\text{on}\,\,{\partial\Omega_{\Gamma}}.$
We will now show actually $\nabla\psi|_{\partial\Omega_{\Gamma}}=0,$ which is automatic when $\Gamma$ is a loop inside $\Omega$. Note that if $\Gamma$ is not a loop in (the interior of) $\Omega$, then its endpoints must lie on $\partial\Omega$ (and may coincide).
@@
If $\Gamma$ touches $\partial\Omega,$ then the portion of $\partial\Omega$ between the two endpoints of $\Gamma$ in $\partial\Omega_{\Gamma}$ lies in $\mathcal{C}(\psi).$
@@
First, we claim that if $\Gamma$ touches $\partial\Omega$, then $d_{\psi}(x)\geq 3$ for all $x\in\Gamma$. Suppose that there is a point $x_{0}\in\Gamma\cap\Omega$ where $\psi$ vanishes at degree $2$. By Lemma 3.1, all the points in $\Gamma$ have to be critical points of $\psi$ with vanishing degree $2$. However, just as in the proof of Corollary 3.2, we have shown that $\Gamma$ cannot touch the boundary. This leads to a contradiction.
Second, $\nabla\psi$ must vanish on $\partial\Omega\cap\partial\Omega_{\Gamma}$. Indeed, take some $x\in\partial\Omega\cap\partial\Omega_{\Gamma}$, if $\nabla\psi(x)\not=0,$ then it follows from the inverse function theorem that $F$ from (3.1) must be analytic in a neighborhood of 0 (as $\psi$ takes zero value in $\partial\Omega$). In addition, the condition that $d_{\psi}(x)\geq 3$ for all $x\in\Gamma$ shows $\psi$ takes the value 0 in $\Gamma$ and that $F(0)=0$. Now as $\Delta\psi=F(\psi)$ with $F$ analytic at 0 and $F(0)=0$, considering the leading term of $F(\psi)$ near $\Gamma$, the vanishing degree of $\Delta\psi$ at $\Gamma$ is greater than or equal to the vanishing degree of $\psi$ at $\Gamma$. This would contradict Lemma 3.3. Hence, the proof of the lemma is complete. ∎
Based on Lemmas 3.4 and 3.2, we have the following corollary concerning $\partial\Omega_{\Gamma}$.
@@
One of the following two cases must hold.
  • Case 1: $\partial\Omega_{\Gamma}$ is an analytic loop in $\Omega$ and contains only critical points of degree 2.
  • Case 2: $\partial\Omega_{\Gamma}$ is a closed loop in $\bar{\Omega}$ containing critical points of degree larger than 2.
Our goal will now be to show that, in both cases, $\psi$ is necessarily a radial function. We will first deal with Case 2, which will be solved essentially by applying existing results from the literature. Case 1 will be divided into two subcases based on whether the function $F$ appeared in (3.1) is analytic at the value of the stream function on $\partial\Omega_{\Gamma}$. These two subcases will be dealt with in the following section using the moving plane method and contradiction argument, respectively.
3.2. Rigidity in Case 2
In the previous subsection, we reduced the problem into studying a semilinear elliptic equation of Serrin type [46]. More precisely, we found a simply connected domain $\Omega_{\Gamma}$ for which $\nabla\psi$ vanishes on $\partial\Omega_{\Gamma}$ and $\Omega_{\Gamma}$ contains only isolated critical points of $\psi$. As a consequence of Lemma 1.3, we deduced that $\psi$ solves (3.1) in $\Omega_{\Gamma}$. In this section, we consider Case 2 of Corollary 3.5, where all points in $\partial\Omega_{\Gamma}$ are critical points with degree greater than or equal to $3$. Our goal is to establish the following proposition.
@@ 3.2. Rigidity in Case 2
Let $\Omega_{\Gamma}$ be a simply connected domain and $\psi$ be an analytic function in $\overline{\Omega_{\Gamma}}$ such that there is no non-isolated critical point of $\psi$ in $\Omega_{\Gamma}\cap\{\psi=0\}$. Now we further assume that there is a continuous function $F$, such that $\psi$ satisfies the following overdetermined problem for an elliptic equation.
(3.2) $\Delta\psi(x)=F(\psi)(x),\text{ for $x\in\Omega_{\Gamma}$,}$
$\psi(x)=\nabla\psi(x)=\nabla^{2}\psi(x)=0,\text{ for all $x\in\partial\Omega_{\Gamma}$}.$
@@
We first recall the following result based on the Steiner symmetrization method.
@@
Let $\Omega$ be an open bounded domain and $\psi$ be a non-negative $C^{2}$ solution to
(3.3) $\left\{\begin{aligned} &\Delta{\psi}(x)=F(\psi)(x),\quad&x\in\Omega,\\ &\psi(x)=0,\ \nabla\psi(x)=0,\nabla^{2}\psi(x)=0,\quad&x\in\partial\Omega.\end{aligned}\right$.
Then $F$ being continuous implies that $\psi$ is locally radial in the sense that
@@
where $A=\cup_{i}A_{i}$, $\mathcal{C}=\{\nabla\psi=0\}$ and $A_{i}$ are disjoint open annuli or balls and $\psi$ is radially symmetric in $A_{i}$.
@@
Given Proposition 3.7, all we need to do is show that $\psi$ is single-signed in $\Omega_{\Gamma}.$ Towards a contradiction, assume that $\psi$ takes both positive and negative values in $\Omega_{\Gamma}.$ Simply by continuity, it follows that the set of zeros of $\psi$ in $\Omega_{\Gamma}$ is infinite. It follows from our choice of $\Gamma$ that at least one of those zeros is not a critical point of $\psi$ (since the critical set in $\Omega_{\Gamma}\cap\{\psi=0\}$ is discrete). It follows that $F$ is analytic in a neighborhood of zero. This contradicts Lemma 3.3 (just as in the proof of Lemma 3.4). ∎
4. Moving Plane Arguments and the Proof of the Main Theorem
We now move on to prove the rigidity in Case 1 of Corollary 3.5, where our basic assumption is that there is no $F$ for which $\psi$ solves (1.4) in all of $\Omega$. More precisely, we may assume that there exists a simply connected domain $\Omega_{\Gamma}$ with $\partial\Omega_{\Gamma}$ analytic and $\psi$ satisfies
(4.1) $\Delta\psi(x)=F(\psi)(x),\text{for $x\in\Omega_{\Gamma}$,}$
@@ 4. Moving Plane Arguments and the Proof of the Main Theorem
$\partial_{nn}\psi=1\quad\text{on}\,\,\partial\Omega_{\Gamma}.$
An important consequence is that $\psi$ increases in the normal direction of the boundary, in a tubular neighborhood of the boundary. In the case where $F$ is $C^{1}$ and $\psi$ is non-negative, we have the following rigidity result due to Serrin [46] (see also the nice review [41]).
Theorem 4.1 (Serrin, [46]).
Assume that $\psi$ satisfies (4.1) and that $\psi>0$ in $\Omega_{\Gamma},$ while $F$ is Lipschitz continuous. Then $\Omega_{\Gamma}$ is a ball and $\psi$ is radial.
@@
One of the following two cases must happen.
  • $F(s)$ is not analytic at $s=0,$ in which case $\psi$ is non-negative in $\Omega_{\Gamma}.$ This is called the unextendable case.
@@
Proof.
To prove the lemma, it suffices to check that $\psi$ is non-negative on $\Omega_{\Gamma}$ when $F$ is not analytic near 0. In fact, the proof is quite similar to that in the proof of Proposition 3.6. Now we assume, towards a contradiction, that there is a point $x_{*}\in\Omega_{\Gamma}$ with $\psi(x_{*})<0$. Since $\nabla\psi=0$ and $\partial_{nn}\psi=1$ on $\partial\Omega_{\Gamma}$, $\psi$ is positive in a tubular neighborhood of $\Gamma$ excluding $\Gamma$ itself. Hence, by continuity, it follows from $\psi(x_{*})<0$ that there are infinitely many zeros of $\psi$ in $\Omega_{\Gamma}$. Since $F$ is not analytic at $s=0,$ the zeros of $\psi$ in $\Omega_{\Gamma}$ must be critical points. In particular, $\psi$ must have infinitely many critical points in $\Omega_{\Gamma}$. This contradicts the choice of $\Gamma$ as the innermost loop. ∎
4.1. The Unextendable Case
This is the case where $F(s)$ is not analytic in $s$ at 0. As mentioned in Lemma 4.2, $\psi$ must be non-negative in $\Omega_{\Gamma}$, since otherwise there would be infinitely many points where $\psi=0$ in $\Omega_{\Gamma}$. Without loss of generality, by scaling and multiplication by a constant factor, we assume that $\max_{\Omega_{\Gamma}}\psi=1$. Analogously as in the proof of Proposition 3.6, the condition that $\psi$ does not have a non-isolated critical point in $\Omega_{\Gamma}$ implies that $F$ is analytic in $(0,1)$. Our goal in this subsection is to prove the following proposition.
@@ 4.1. The Unextendable Case
Let $\Omega_{\Gamma}$ be a simply connected analytic domain, and $\psi$ be a non-negative analytic function in $\overline{\Omega_{\Gamma}}$. Assume $F\in C([0,1])$ is analytic on $(0,1).$ If $F$ is not analytic at 0 and $\psi$ satisfies the over-determined problem (4.1) for elliptic equation, then $\Omega_{\Gamma}$ is a ball and $\psi$ is radial.
@@
Without loss of generality, we may assume $F(0)=1,$ since the case $F(0)=0$ has already been covered by Proposition 3.6.
For the proof of Proposition 4.3, it suffices to prove that $\psi$ is even symmetric to all directions $e\in\mathbb{S}^{1}$. Without loss of generality, by rotation, we need only to prove the symmetry with respect to $e=(1,0)$. The key is to use the moving plane method to show that $\psi$ is even symmetric to $(1,0)$. We start with a series of lemmas concerning the regularity of $F$ and the critical set of $\psi$. First, either $F$ is analytic at $1$ or $F$ is not smooth and admits a Puiseux expression near $1$.
@@
When the maxima of $\psi$ are critical points of degree $2k_{0},$ for some $k_{0}\geq 1,$ $F(s)$ admits the following convergent expansion, for $0\leq 1-s\ll 1$,
(4.2) $F(s)=\sum_{k=k_{0}-1}^{\infty}a_{k}(1-s)^{\frac{k}{k_{0}}},\text{ with $a_{k}$ real and $a_{k_{0}-1}<0$. }$
@@
Fix a point $x^{0}\in\Omega_{\Gamma}$ such that $\psi(x^{0})=1$, as $\psi$ attains the global maximum at $x^{0}$, we have $d_{\psi}(x^{0})$ is an even number. In the case $d_{\psi}(x^{0})=2$, without loss of generality, by a suitable rotation and translation, we assume $x^{0}=0$ and $\partial_{11}\psi(0)<0$. As $\partial_{1}\psi(0)=0$ and $\partial_{11}\psi(0)<0$, for $0\leq 1-s\ll 1$, the function $\psi(t,0)$ near 0 has exactly two inverses $f_{+}(s)$ and $f_{-}(s)$, with $f_{+}(s)\geq 0$ and $f_{-}(s)\leq 0$.
We show that $f_{+}$ and $f_{-}$ admit a Puisuex expansion near $1$. In particular, we will show that there is an analytic function $G$, such that $f_{+}(s)=G(\sqrt{1-s})$ and $f_{-}(s)=G(-\sqrt{1-s})$. Indeed, given the Taylor expression of $\psi(t,0)=1+a_{2}t^{2}+\sum_{n\geq 3}a_{n}t^{n}$ with $a_{2}<0$, we have
(4.3) $\psi(f_{\pm}(s),0)=s\Leftrightarrow$ $f_{\pm}^{2}(s)\left(1+\sum_{k\geq 3}\frac{a_{k}}{a_{2}}f_{\pm}^{k-2}(s)\right)=\frac{s-1}{a_{2}}$
$\Leftrightarrow$ $f_{\pm}(s)\sqrt{1+\sum_{k\geq 3}\frac{a_{k}}{a_{2}}f_{\pm}^{k-2}(s)}=\pm\sqrt{\frac{s-1}{a_{2}}}.$
Note that the function $s\sqrt{1+\sum_{k\geq 3}\frac{a_{3}}{a_{2}}s^{k-2}}$ is an analytic function with a non-vanishing derivative at 0, hence it admits an analytic inverse $H$. Therefore, we can correspondingly choose $G(s)=H(\frac{s}{\sqrt{-a_{2}}})$ so that $f_{\pm}(s)=G(\pm\sqrt{1-s})$.
With the aid of $_{+}, ff_{-}$, and $G$, for $0\leq 1-s\ll 1$, we have
@@
The fact that $\Delta\psi(G(\sqrt{1-s}),0)$ agrees with $\Delta\psi(G(-\sqrt{1-s}),0)$, implies that all the terms consisting of odd powers of $\sqrt{1-s}$ vanish in the Puiseux expansion of $F$ near $1$. Therefore, $F$ is analytic near $1$.
Now in the case where $d_{\psi}(x^{0})=2k_{0}\geq 4$, again assume that $x^{0}=0$ and $\partial_{1}^{2k_{0}}\psi<0$, for $0\leq 1-s\ll 1$, the function $\psi(t,0)$ near 0 has exactly two inverses $f_{+}(s)$, $f_{-}(s)$. Similarly, as in the case where $d_{\psi}(x^{0})=2$, it follows that $F$ near $1$ has the Puiseux expansion. More precisely, for $0\leq 1-s\ll 1$, we have (4.2). This finishes the proof of the lemma. ∎
Analogously, by taking a normal line segment through a boundary point $x\in\partial\Omega_{\Gamma}$, we have $F(s)$ is analytic in the variable $\sqrt{s}$ near 0. Together with the condition that $F$ is not analytic near 0, we have the following Puiseux expansion of $F$ near 0.
@@
$F$ near 0 admits a convergent Puiseux series expression: for $0\leq s\ll 1$, one has
(4.4) $F(s)=\sum_{k=0}^{\infty}a_{k}s^{\frac{k}{2}},$
where $a_{k}\in\mathbb{R}$, $a_{0}=1$ and $a_{k_{0}}\neq 0$ for an odd number $k_{0}$.
Moreover, arguing similarly as in the case $d_{\psi}(x^{0})=2$ in the proof of Lemma 4.4, the condition $F$ is not analytic at 0 implies that $\psi$ is positive in $\Omega_{\Gamma}$. Together with the fact
@@
we can show that $\psi$ has a uniform lower bound in $\Omega_{\Gamma}$.